# Bootstrap3D: Improving 3D Content Creation with Synthetic Data

### v1

@@ Abstract
Recent years have witnessed remarkable progress in multi-view diffusion models for 3D content creation. However, there remains a significant gap in image quality and prompt-following ability compared to 2D diffusion models. A critical bottleneck is the scarcity of high-quality 3D assets with detailed captions. To address this challenge, we propose Bootstrap3D, a novel framework that automatically generates an arbitrary quantity of multi-view images to assist in training multi-view diffusion models. Specifically, we introduce a data generation pipeline that employs (1) 2D and video diffusion models to generate multi-view images based on constructed text prompts, and (2) our fine-tuned 3D-aware MV-LLaVA for filtering high-quality data and rewriting inaccurate captions. Leveraging this pipeline, we have generated 1 million high-quality synthetic multi-view images with dense descriptive captions to address the shortage of high-quality 3D data. Furthermore, we present a Training Timestep Reschedule (TTR) strategy that leverages the denoising process to learn multi-view consistency while maintaining the original 2D diffusion prior. Extensive experiments demonstrate that Bootstrap3D can generate high-quality multi-view images with superior aesthetic quality, image-text alignment, and maintained view consistency.
@@ 1 Introduction
3D content creation stands as a fundamental challenge within the generative domain, boasting widespread applications in augmented reality (AR) and game modeling. Unlike 2D image generation, the dearth of high-quality 3D models persists as a significant hurdle to overcome. In the realm of 2D image generation, the pivotal role of training on billion-scale image-text pairs [64] has been firmly established [5, 62, 40, 12, 11]. However, in 3D content generation, the scarcity of high-quality 3D models often compels reliance on the priors of 2D diffusion models. The predominant methodologies in this domain can be categorized into two main streams: 1) Gaining optimized neural representations from fixed 2D diffusion models via Score Distillation Sampling (SDS) loss [59, 67, 49, 66, 47, 80], which are time-intensive, lacking diversity and suffer from low robustness although capable of producing high-quality 3D objects. 2) Fine-tuning 2D diffusion models to achieve multi-view generation [41, 66, 67] , directly synthesizing 3D objects through sparse reconstruction models [41, 79, 88, 89, 70, 84]. With recent improvements in large-scale sparse view reconstruction models and 3D representations [36], the second stream is garnering increasing attention.
@@
Fine-tuning 2D diffusion models for multi-view generation remains challenging owing to the insufficiency in both data quality and quantity. Previous methods [60, 41, 67, 19] primarily train on a filtered subset of high-quality data from Objaverse [20] and Objaverse-XL [19]. The scarcity of high-quality data often introduces various shortcomings. In single-view based novel view synthesis [49, 66, 78, 76], if the input images deviate from the distribution of the training data, it can induce issues such as motion blurring, object distortion and deformation [66]. Moreover, in direct text-to-multi-view image generation, the pursuit of enhancing view consistency compromises the aesthetic and photo-realistic quality. For instance, Intant3D [41] fine-tunes SDXL [58] using only 10K high-quality Objaverse [20] data with a small learning rate for 10K steps, which does not fundamentally prevent the catastrophic forgetting problem of losing 2D diffusion prior, leading to compromised image quality. Recent endeavors have predominantly focused on alleviating data scarcity and improving view consistency from a model-centric perspective [35, 66, 72], with limited exploration into the improvement of training data itself.
@@
Recent Multimodal Large Language Models (MLLMs) [46, 13, 42, 2, 3] like GPT-4V [55] and Gemini [73], possess image understanding capabilities and rudimentary 3D world awareness, has enabled automatic quality assessment of multi-view images and dense caption generation. Furthermore, notable advancements in video diffusion [7, 76] have improved the generalizability of novel view synthesis [76, 15, 39]. Employing these advancements, we propose Bootstrap3D to generate synthetic data to counteract the data deficiencies inherent in training multi-view diffusion models. To be specific, we introduce the Bootstrap3D data generation pipeline for producing high-quality multi-view images with dense descriptive captions. Subsequently, we fine-tune a multi-view-aware MLLM model, dubbed as MV-LLaVA. MV-LLaVA achieves fully automated high-quality data generation with both efficiency and accuracy. To mitigate catastrophic forgetting of 2D diffusion prior, we introduce a training timestep reschedule (TTR) strategy when fine-tuning multi-view diffusion models. Specifically, we use the phased nature of the denoising process [30] and limit different training time steps for synthetic data and real data to achieve enhanced image quality with maintained view consistency.
@@
Through extensive experiments, we demonstrate that our method significantly enhances the adherence of the multi-view diffusion model to text prompts and image quality while ensuring view consistency. Integrated with the reconstruction model and Gaussian representations, our approach facilitates the creation of 3D models with superior quality.
Our contributions are summarized into the following points:
@@
1) We present an automated Bootstrap3D data generation pipeline that uses the video diffusion model and our fine-tuned 3D-aware MV-LLaVA to synthesize an arbitrary number of high-quality 3D data.
@@ 2 Related Work
Existing 3D datasets and data pre-processing. Existing object level 3D datasets, sourced either from CAD [10, 87, 20, 19] or scan from real objects [1, 91, 22, 86], are still small in size. Most state-of-the-art open-sourced 3D content creation models are trained on Objaverse [20] (800k) and Objaverse-XL [19] (10M). However, there still exists a huge gap compared to data used for training 2D diffusion models like Laion5B [64]. In addition to quantity, quality is also an important problem remains to be solved as many methods [19, 67, 41, 60, 70, 72, 89] trained on Objaverse or Objaverse-XL rely on simple methods like CLIP [61] score to filter out low-quality data, making the precious 3D data even less. This gap poses a significant obstacle to achieving high-quality 3D content creation compared to its 2D diffusion counterpart. Another critical gap that requires attention is the quality of the 3D object’s caption. Previous work Cap3D [52] propose to apply BLIP-2 [42] and GPT-4 [56] to generate caption based on multi-view images. However, this approach, without direct input image into GPT, can lead to severe hallucination. Given recent breakthroughs in improving text-image alignment through caption rewriting [5, 12, 11, 24], there is a pressing need to rewrite denser and more accurate captions for 3D objects with the assistance of advanced Multimodal Large Language Models (MLLMs) compared to what Cap3D [52] have accomplished. In this work, we propose a new data generation pipeline to synthesize multi-view images and rewrite captions for 3D objects incorporating additional quality scoring mechanisms to address the aforementioned issues.
@@
Text-to-3D content creation. The field of 3D content creation has been a vibrant area of research over the past years. One prominent research direction explores the use of Score Distillation Sampling (SDS) [59] and its variants [14, 18, 28, 44, 45, 49, 67, 50, 51, 80, 71, 77, 90], using the priors of 2D diffusion models to optimize 3D representations. While these methods have demonstrated success in producing high-quality 3D generations, they often require prolonged optimization time to converge. In contrast, recent studies [32, 79, 41, 70, 74, 89, 84] have proposed the direct inference of 3D representations [54, 9, 36, 92] conditioned by images. Among these approaches, Instant3D [41] stands out by utilizing multi-view images of the same object to directly deduce the Triplane [9] representation. This approach effectively addresses the issue of ambiguous unseen areas inherent in the single image to 3D conversions, as encountered in LRM [32] and TripoSR [74]. Instant3D, along with subsequent works [89, 94, 81, 88], efficiently decomposes 3D generation into two processes: the generation of multi-view images using multi-view diffusion model [49, 50, 47, 67, 48, 66, 51, 35, 76] and large reconstruction model to generate 3D representations conditioned on these multi-view images. In this work, we introduce a method that significantly enhances the scalability of training and data generation for multi-view image generation.
@@
Video diffusion for novel view synthesis. Recent advancements in video diffusion have marked a significant breakthrough, with models such as Sora [7] and SVD [6] scaling up the direct generation process from images to videos. Following these developments, a series of works [82, 39, 6, 53, 27, 15] represented by SV3D [76], have fine-tuned these video diffusion models for 3D content creation. Despite these groundbreaking developments, the new perspective images generated based on video priors still suffer from issues like motion blur, leading to inconsistencies with natural images. In this work, we propose to utilize SV3D [76] as a data generator to produce novel views of given images with additional quality checks to leave only high-quality data.
@@
Multimodal Large Language Models. With the development of large language models [8, 56, 16, 3, 31, 75], multimodal large language models (MLLMs) [93, 2, 42, 43, 33, 23, 4, 46, 21, 69], such as GPT-4V [55], have demonstrated groundbreaking 2D comprehension capabilities and open-world knowledge. As is discovered in GPTEval3D [85], GPT-4V can achieve human-aligned evaluation for multi-view images rendered from 3D objects. In this work, we fine-tune the LLaVA [46], a lightweight MLLM for quality judgment and descriptive caption generation based on multi-view images. This integration uses the strengths of MLLMs in understanding visual content to enhance the quality assessment and captioning process, contributing to improved 3D content generation.
@@ 3 Methods
Due to the scarcity of high-quality 3D data, we develop the Bootstrap3D data generation pipeline to efficiently construct an arbitrary number of training data (Sec. 3.1). Subsequently, the quality of generated multi-view images is assessed using the powerful GPT-4V [55] or our proposed MV-LLaVA [46] model to generate dense descriptive captions efficiency and faithfully (Sec. 3.2). We also design a training timestep reschedule (Sec. 3.3) when fine-tuning the multi-view diffusion model with our synthetic and real data.
@@ 3.1 Bootstrap3D Data Generation Pipeline
As illustrated in Fig.1, our data generation pipeline initially employs GPT-4 [55] to generate a multitude of imaginative and varied text prompts [85]. Subsequently, to generate 2D images that closely align with the text prompts, we utilize the PixArt-Alpha [12] model use FlanT5 [17] text encoder with DiT [57] architecture for text-to-image (T2I) generation. Thereafter, we use SV3D [76] for novel view synthesis. Given the significant motion blur and distortion often present in SV3D [76] outputs, we further employ Multimodal Large Language Models(MLLM) to evaluate the quality of multi-view images. To rectify mismatches between multi-view images and the original text prompts induced by novel view synthesis and provide more precise captions, we further propose MV-LLaVA to generate dense descriptive captions for multi-view images.
@@ 3.2 Multi-View LLaVA (MV-LLaVA)
To efficiently generate captions and label quality scores for both generated multi-view images and 3D assets in Objaverse [20], we propose the Multi-View LLaVA (MV-LLaVA) that fine-tune LLaVA [46] based on our instructive conversation pairs generated by the powerful GPT-4V [55].
Preparing the instruction tuning data. As shown in Fig.1, we use GPT-4 to generate 20k varied text prompts based on prompts designed in  [85] and use PixArt-alpha [12] to generate single view image and use SV3D [76] or Zero123++ [66] to generate multi-view images. For this 20k generated multi-view images, we prompt GPT-4V [55] to generate comments on view consistency, image quality and generate dense descriptive captions. For the additional 10K rendered multi-view images from Objaverse [20], we prompt GPT-4V (detailed prompts in Sup. D.1) to offer feedback on the quality and aesthetic appeal of 3D objects, along with style judgments. We utilize these 30K high-quality multi-view image text pairs (prompts detailed in Sup. D.2) as the instruction tuning data for LLaVA.
Figure 2: MV-LLaVA. We use GPT-4V [55] to generate long descriptive captions, quality scoring, and reasoning processes for multi-view images to construct the instruction tuning dataset. Then we fine-tune our MV-LLaVA based on LLaVA [46] to serve as the human-aligned quality checker and captioner for multi-view images.
Instruction tuning. As presented in the left part of Fig. 2, due to the LLaVA’s maximum training context length constraints of 2048, we input four images separately into CLIP-L/14 [61] and generate 4$\times$256 image tokens. Inspired by ShareGPT-4V [13], we freeze only a portion of layers of CLIP [61] in the first stage of pre-training to enhance multi-view awareness and texture perception of vision encoder (detailed in Sup. C.1). As shown in the right part of Fig. 2, we first ask the model to generate descriptions, then let the model score the quality based on multi-view images and captions. Our approach encourages LLM to deduct more reasonable quality scores through chain-of-thought [83]. A mixture of original training data of LLaVA is adopted to mitigate over-fitting. As a result, we obtain MV-LLaVA, which efficiently filters and re-captions both synthetic data and 3D assets. As detailed in Sup.C, MV-LLaVA can not only generate more accurate, less hallucinated dense captions that faithfully describe 3D objects compared to Cap3D [52] but also assign the human-aligned quality score on both synthetic data and Objaverse assets. The filtered high-quality multi-view images with rewritten dense captions served as training data for the diffusion model.

### v2

# Bootstrap3D: Improving Multi-view Diffusion Model with Synthetic Data

@@ Abstract
Recent years have witnessed remarkable progress in multi-view diffusion models for 3D content creation. However, there remains a significant gap in image quality and prompt-following ability compared to 2D diffusion models. A critical bottleneck is the scarcity of high-quality 3D data with detailed captions. To address this challenge, we propose Bootstrap3D, a novel framework that automatically generates an arbitrary quantity of multi-view images to assist in training multi-view diffusion models. Specifically, we introduce a data generation pipeline that employs (1) 2D and video diffusion models to generate multi-view images based on constructed text prompts, and (2) our fine-tuned 3D-aware MV-LLaVA for filtering high-quality data and rewriting inaccurate captions. Leveraging this pipeline, we have generated 1 million high-quality synthetic multi-view images with dense descriptive captions to address the shortage of high-quality 3D data. Furthermore, we present a Training Timestep Reschedule (TTR) strategy that leverages the denoising process to learn multi-view consistency while maintaining the original 2D diffusion prior. Extensive experiments demonstrate that Bootstrap3D can generate high-quality multi-view images with superior aesthetic quality, image-text alignment, and maintained view consistency.
@@ 1 Introduction
3D content creation stands as a fundamental challenge within the generative domain, boasting widespread applications in augmented reality (AR) and game modeling. Unlike 2D image generation, the dearth of high-quality 3D models persists as a significant hurdle to overcome. In the realm of 2D image generation, the pivotal role of training on billion-scale image-text pairs (Schuhmann et al., 2022) has been firmly established (Betker et al., 2023; Rombach et al., 2022; Li et al., 2024; Chen et al., 2023a; 2024a). However, in 3D content generation, the scarcity of high-quality 3D models often compels reliance on the priors of 2D diffusion models. The predominant methodologies in this domain can be categorized into two main streams: 1) Gaining optimized neural representations from fixed 2D diffusion models via Score Distillation Sampling (SDS) loss (Poole et al., 2022; Shi et al., 2023b; Liu et al., 2023b; Shi et al., 2023a; Liu et al., 2023a; Wang et al., 2024a), which are time-intensive, lacking diversity and suffer from low robustness although capable of producing high-quality 3D objects. 2) Fine-tuning 2D diffusion models to achieve multi-view generation (Li et al., 2023a; Shi et al., 2023a; b) , directly synthesizing 3D objects through sparse reconstruction models (Li et al., 2023a; Wang et al., 2023b; Xu et al., 2024a; b; Tang et al., 2024a; Wei et al., 2024). With recent improvements in large-scale sparse view reconstruction models and 3D representations (Kerbl et al., 2023), the second stream is garnering increasing attention.
@@
Fine-tuning 2D diffusion models for multi-view generation remains challenging owing to the insufficiency in both data quality and quantity. Previous methods (Qiu et al., 2023; Li et al., 2023a; Shi et al., 2023b; Deitke et al., 2024) primarily train on a filtered subset of high-quality data from Objaverse (Deitke et al., 2023) and Objaverse-XL (Deitke et al., 2024). The scarcity of high-quality data often introduces various shortcomings. In single-view based novel view synthesis (Liu et al., 2023b; Shi et al., 2023a; Wang & Shi, 2023; Voleti et al., 2024), if the input images deviate from the distribution of the training data, it can induce issues such as motion blurring, object distortion and deformation (Shi et al., 2023a).
Figure 1: Bootstrap3D can generate high quality multi-view images with precise long text control and style customization while maintaining view consistency.
Moreover, in direct text-to-multi-view image generation, the pursuit of enhancing view consistency compromises the aesthetic and photo-realistic quality. For instance, Intant3D (Li et al., 2023a) fine-tunes SDXL (Podell et al., 2023) using only 10K high-quality Objaverse (Deitke et al., 2023) data with a small learning rate for 10K steps, which does not fundamentally prevent the catastrophic forgetting problem of losing 2D diffusion prior, leading to compromised image quality. Recent endeavors have predominantly focused on alleviating data scarcity and improving view consistency from a model-centric perspective (Kant et al., 2024; Shi et al., 2023a; Tang et al., 2024b), with limited exploration into the improvement of training data itself.
@@
Recent Multimodal Large Language Models (MLLMs) (Liu et al., 2024a; Chen et al., 2023b; Li et al., 2023b; Alayrac et al., 2022; Anil et al., 2023) like GPT-4V (OpenAI, 2023a) and Gemini (Team et al., 2023), possess image understanding capabilities and rudimentary 3D world awareness, has enabled automatic quality assessment of multi-view images and dense caption generation. Furthermore, notable advancements in video diffusion (Brooks et al., 2024; Voleti et al., 2024) have improved the generalizability of novel view synthesis (Voleti et al., 2024; Chen et al., 2024b; Kwak et al., 2023). Employing these advancements, we propose Bootstrap3D to generate synthetic data to counteract the data deficiencies inherent in training multi-view diffusion models. To be specific, we introduce the Bootstrap3D data generation pipeline for producing high-quality multi-view images with dense descriptive captions. Subsequently, we fine-tune a multi-view-aware MLLM model, dubbed as MV-LLaVA, to achieve fully automated high-quality data annotation with both efficiency and accuracy. To mitigate catastrophic forgetting of 2D diffusion prior, we introduce a training timestep reschedule (TTR) strategy when fine-tuning multi-view diffusion models. Specifically, we use the phased nature of the denoising process (Ho et al., 2020) and limit different training time steps for synthetic data to achieve enhanced image quality with maintained view consistency.
@@
Through extensive experiments, we demonstrate that our method significantly enhances the adherence of the multi-view diffusion model to text prompts and image quality while ensuring view consistency. Integrated with the reconstruction model, our approach facilitates the creation of 3D models with superior quality. Our contributions are summarized into the following points:
@@
1) We present an automated Bootstrap3D data generation pipeline that uses the video diffusion model and our fine-tuned 3D-aware MV-LLaVA to synthesize an arbitrary number of high-quality multi-view image text pairs.
@@ 2 Related Work
Existing 3D datasets and data pre-processing. Existing object level 3D datasets, sourced either from CAD (Chang et al., 2015; Wu et al., 2015; Deitke et al., 2023; 2024) or scan from real objects (Aanæs et al., 2016; Yao et al., 2020; Downs et al., 2022; Wu et al., 2023), are still small in size. Most state-of-the-art open-sourced 3D content creation models are trained on Objaverse (Deitke et al., 2023). However, there still exists a huge gap compared to data used for training 2D diffusion models (Schuhmann et al., 2022). In addition to quantity, quality is also an important problem remains to be solved as many methods (Shi et al., 2023b; Li et al., 2023a; Qiu et al., 2023; Tang et al., 2024a) trained on Objaverse rely on filtering out low-quality data, making the precious 3D data even less. Another critical gap that requires attention is the quality of the 3D object’s caption. Previous work Cap3D (Luo et al., 2024) propose to apply BLIP-2 (Li et al., 2023b) and GPT-4 (OpenAI, 2023b) to generate caption based on multi-view images. However, this approach, without direct input image into GPT, can lead to severe hallucination. Given recent breakthroughs in improving text-image alignment through caption rewriting (Betker et al., 2023; Chen et al., 2023a; 2024a; Esser et al., 2024), there is a pressing need to rewrite denser and more accurate captions for 3D objects with the assistance of advanced Multimodal Large Language Models (MLLMs). In this work, we propose a new data generation pipeline to synthesize multi-view images and rewrite captions for 3D objects incorporating additional quality scoring mechanisms to address the aforementioned issues.
@@
Text-to-3D content creation. The field of 3D content creation has been a vibrant area of research over the past years. One prominent research direction explores the use of Score Distillation Sampling (SDS) (Poole et al., 2022) and its variants (Chen et al., 2023c; Chung et al., 2023; Hertz et al., 2023; Liang et al., 2023; Lin et al., 2023; Liu et al., 2023b; Shi et al., 2023b; Liu et al., 2023c; Long et al., 2023; Wang et al., 2024a; Tang et al., 2023; Wang et al., 2023a; Yang et al., 2024; Qi et al., 2024), using the priors of 2D diffusion models to optimize 3D representations. While these methods have demonstrated success in producing high-quality 3D generations, they often require prolonged optimization time to converge. In contrast, recent studies (Hong et al., 2023; Wang et al., 2023b; Li et al., 2023a; Tang et al., 2024a; Tochilkin et al., 2024; Xu et al., 2024b; Wei et al., 2024) have proposed the direct inference of 3D representations (Mildenhall et al., 2021; Chan et al., 2022; Kerbl et al., 2023; Zhang et al., 2023a) conditioned by images. Among these approaches, Instant3D (Li et al., 2023a) stands out by utilizing multi-view images of the same object to directly deduce the Triplane (Chan et al., 2022) representation. This approach effectively addresses the issue of ambiguous unseen areas inherent in the single image to 3D conversions, as encountered in LRM (Hong et al., 2023) and TripoSR (Tochilkin et al., 2024). Instant3D, along with subsequent works (Xu et al., 2024b; Zheng et al., 2024; Wang et al., 2024b; Xu et al., 2024a), efficiently decomposes 3D generation into two processes: the generation of multi-view images using multi-view diffusion model (Liu et al., 2023b; c; a; Shi et al., 2023b; Liu et al., 2024b; Shi et al., 2023a; Long et al., 2023; Kant et al., 2024; Voleti et al., 2024) and large reconstruction model to generate 3D representations conditioned on these multi-view images. In this work, we introduce a method that significantly enhances the scalability of training and data generation for multi-view image generation.
@@
Video diffusion for novel view synthesis. Recent advancements in video diffusion have marked a significant breakthrough, with models such as Sora (Brooks et al., 2024) and SVD (Blattmann et al., 2023) scaling up the direct generation process from images to videos. Following these developments, a series of works (Wang et al., 2023c; Kwak et al., 2023; Blattmann et al., 2023; Melas-Kyriazi et al., 2024; Han et al., 2024; Chen et al., 2024b) represented by SV3D (Voleti et al., 2024), have fine-tuned these video diffusion models for 3D content creation. Despite these groundbreaking developments, the new perspective images generated based on video priors still suffer from issues like motion blur. In this work, we propose to utilize SV3D (Voleti et al., 2024) as a data generator to produce novel views of given images with additional quality checks to leave only high-quality data.
@@
Multimodal Large Language Models. With the development of large language models (Brown et al., 2020; OpenAI, 2023b; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Touvron et al., 2023), multimodal large language models (MLLMs) (Zhang et al., 2023b; Alayrac et al., 2022; Li et al., 2023b; 2022; Huang et al., 2023; Driess et al., 2023; Awadalla et al., 2023; Liu et al., 2024a; Dong et al., 2024; Sun et al., 2023), such as GPT-4V (OpenAI, 2023a), have demonstrated groundbreaking 2D comprehension capabilities and open-world knowledge. As is discovered in GPTEval3D (Wu et al., 2024), GPT-4V can achieve human-aligned evaluation for multi-view images rendered from 3D objects. In this work, we fine-tune the LLaVA (Liu et al., 2024a) for quality judgment and descriptive caption generation based on multi-view images.
@@ 3 Methods
Due to the scarcity of high-quality 3D data, we develop the Bootstrap3D data generation pipeline to efficiently construct an arbitrary number of training data (Sec. 3.1). Subsequently, the quality of generated multi-view images is assessed using the powerful GPT-4V (OpenAI, 2023a) or our proposed MV-LLaVA (Liu et al., 2024a) model to generate dense descriptive captions efficiency and faithfully (Sec. 3.2). We also design a training timestep reschedule (Sec. 3.3) when fine-tuning the multi-view diffusion model with our synthetic and real data.
@@ 3.1 Bootstrap3D Data Generation Pipeline
As illustrated in Fig.2, our data generation pipeline initially employs GPT-4 (OpenAI, 2023a) to generate a multitude of imaginative and varied text prompts (Wu et al., 2024). Subsequently, to generate 2D images that closely align with the text prompts, we utilize the PixArt-Alpha (Chen et al., 2023a) model use FlanT5 (Chung et al., 2024) text encoder with DiT (Peebles & Xie, 2023) architecture for text-to-image (T2I) generation. Thereafter, we use SV3D (Voleti et al., 2024) for novel view synthesis. Given the significant motion blur and distortion often present in SV3D (Voleti et al., 2024) outputs, we further employ Multimodal Large Language Models(MLLM) to evaluate the quality of multi-view images. To rectify mismatches between multi-view images and the original text prompts induced by novel view synthesis and provide more precise captions, we further propose MV-LLaVA to generate dense descriptive captions for multi-view images.
@@ 3.2 Multi-View LLaVA (MV-LLaVA)
To efficiently generate captions and label quality scores for both generated multi-view images and 3D assets in Objaverse (Deitke et al., 2023), we propose the Multi-View LLaVA (MV-LLaVA) that fine-tune LLaVA (Liu et al., 2024a) based on our instructive conversation pairs generated by the powerful GPT-4V (OpenAI, 2023a).
Preparing the instruction tuning data. As shown in Fig.2, we use GPT-4 to generate 20k varied text prompts based on prompts designed in (Wu et al., 2024) and use PixArt-alpha (Chen et al., 2023a) to generate single view image and use SV3D (Voleti et al., 2024) or Zero123++ (Shi et al., 2023a) to generate multi-view images. For these 20k generated multi-view images, we prompt GPT-4V (OpenAI, 2023a) to generate comments on view consistency, image quality and generate dense descriptive captions. For the additional 10K rendered multi-view images from Objaverse (Deitke et al., 2023), we prompt GPT-4V (detailed prompts in Sup. A.5.1) to offer feedback on the quality and aesthetic appeal of 3D objects, along with style judgments. We utilize these 30K high-quality multi-view image text pairs (prompts detailed in Sup. A.5.2) as the instruction tuning data for LLaVA.
Figure 3: MV-LLaVA. We use GPT-4V (OpenAI, 2023a) to generate long descriptive captions, quality scoring, and reasoning processes for multi-view images to construct the instruction tuning dataset. Then we fine-tune our MV-LLaVA based on LLaVA (Liu et al., 2024a) to serve as the human-aligned quality checker and captioner for multi-view images.
Instruction tuning. As presented in the left part of Fig. 3, due to the LLaVA’s maximum training context length constraints of 2048, we input four images separately into CLIP-L/14 (Radford et al., 2021) and generate 4$\times$256 image tokens. Inspired by ShareGPT-4V (Chen et al., 2023b), we freeze only a portion of layers of CLIP (Radford et al., 2021) in the first stage of pre-training to enhance multi-view awareness and texture perception of vision encoder (detailed in Sup. A.4.1). As shown in the right part of Fig. 3, we first ask the model to generate descriptions, then let the model score the quality based on multi-view images and captions. Our approach encourages LLM to deduct more reasonable quality scores through chain-of-thought (Wei et al., 2022). A mixture of original training data of LLaVA is adopted to mitigate over-fitting. As a result, we obtain MV-LLaVA, which efficiently filters and re-captions both synthetic data and 3D assets. As detailed in Sup.A.4, MV-LLaVA can not only generate more accurate, less hallucinated dense captions that faithfully describe 3D objects compared to Cap3D (Luo et al., 2024) but also assign the human-aligned quality score on both synthetic data and Objaverse assets. The filtered high-quality multi-view images with rewritten dense captions served as training data for the diffusion model.